IS is Y = 2500-50i, and the interest rate reaction function is ip = 2 + 0.5(P-P*), where ip is the central bank policy rate. Assume that i (interest rate) is 0.5 + ip (interest rate is a wedge over the policy rate)
1
Expert's answer
2020-05-17T17:58:22-0400

"i=0.5+2+0.5(P-P^*)=2.5+0.5(P-P^*)"

"Y=2500+50(2.5+0.5(P-P^*))=2500+125+25(P-P^*)=2625+25(P-P^*)"

IS is determined by function Y. P-price in the current period P*-price in the base period

IS depends on the change of price